# Characterization of Silicon Photomultiplier Photon Detection Efficiency at Liquid Nitrogen Temperature

### v1

@@ 4 Results
The PDE was measured at room temperature and at liquid nitrogen temperature for the KETEK SiPM and Hamamatsu SiPM, see Fig. 3. Errors on all three quantities in Eq. 2.2 were propagated during error calculations. The room temperature was 72 F (295 K) held by an air conditioning unit and the liquid nitrogen temperature was measured to be 77 K with a calibrated PT1000 sensor placed in between the SiPM board and copper cold finger. These temperatures were stable within 1 K during data taking. The breakdown voltage of the KETEK SiPM was measured at room temperature to be 24.1 V; the breakdown voltage at 77 K was 20.6 V. The breakdown voltages were obtained by extrapolating the SiPM gain, $G$, as a function of reverse bias to $G=1$ and were confirmed with a reverse IV curve from a Keithley 2450 Sourcemeter®. Likewise, the breakdown voltage of the Hamamatsu SiPM was measured at room temperature to be 51.5 V; the breakdown voltage at 77 K was 42.2 V. Because the breakdown voltage depends on the temperature, the breakdown voltage was also used as a cross-check on the temperature. The breakdown voltages at liquid nitrogen temperature were determined by submerging the devices completely in liquid nitrogen in a separate experiment prior to the PDE measurement.
Figure 3: Measurement of the PDE for a KETEK PM3325-WB-D0 SiPM and a Hamamatsu S13360-3050CS SiPM for a range of overvoltages. Measurements are done at room temperature and liquid nitrogen temperature using $562.5$ nm light. The datasheet values for the PDEs at the peak LED wavelength used are provided for reference.
The PDE was measured for a range of overvoltages and results show a significant decrease in PDE at liquid nitrogen temperature for all overvoltages. This decrease in PDE is present even when accounting for possible changes in the optical properties of the setup at different temperatures or between configurations. Any systematic impact of the light uniformity in the integrating sphere or SiPM alignment would be seen during room temperature absolute PDE measurements. Our room temperature PDE measurements are in good agreement with the manufacturer’s datasheets, as seen in Fig. 3. Systematic studies were performed to determine if there was any temperature dependence in the transmission coefficient of the optical fibers connecting to the reference SiPM, the emission spectrum of the LED, or the reflectivity of the integrating sphere.
@@
Figure 4: The fluctuation in the effective PDE, $\frac{N_{\mathrm{SiPM}}^{\mathrm{rel.}}}{N_{\mathrm{ref.}}^{\mathrm{rel.}}}$, when only the optical fibers were cooled down sets a limit on the systematic error at $2\%$.
We also investigated the systematics arising from the finite width of the LED’s emission spectrum. Using Eqn. 2.3 to calculate the reference SiPM PDE instead of using the reference SiPM’s PDE at the peak emission wavelength of the LED introduces only an average $0.02\%$ difference in our final cryogenic PDE results, which is below our statistical sensitivity. We therefore compute the cryogenic PDE using only the peak emission wavelength of the LED.
We checked that $\rho(\lambda)$ (the LED’s emission spectrum) and the integrating sphere’s optical properties were not functions of temperature. To do so, we compared the number of photons detected at the reference SiPM at 77 K and 295 K while the test stand was in the relative PDE measurement configuration. The reference SiPM has a comparable spectral response to the SiPMs under test. The results, shown in Fig. 5, indicate that the reference SiPM detects roughly the same number of photons when the test stand is at both room and cryogenic temperatures. Thus there is neither a shift in the emission spectrum of the LED nor a shift in the reflectivity of the integrating sphere. This measurement also serves as an in situ cross-check that the optical fibers’ transmission probability is not temperature dependent. Therefore, any changes in the optical properties of the test stand between the two temperatures do not account for the observed decrease in PDE.
@@
The decrease in PDE at 77 K indicates that some combination of the quantum efficiency, avalanche triggering probability, and effective fill factor may be changing with temperature. Similar reductions of the PDE at cryogenic temperatures have been found in other SiPMs at 10 K using a calibrated system [13], and at 77 K using uncalibrated systems [14], [15] and calibrated systems [16]. Since silicon is an indirect band-gap semiconductor and fewer phonons are present in the crystal lattice at liquid nitrogen temperature, the probability of charge carriers moving to the conduction band from the absorption of a photon will be lower. This is reflected in the increase of the absorption depth of photons at a given energy at liquid nitrogen temperatures. The band-gap of silicon increases with decreasing temperature, but this may produce negligible effects on the quantum efficiency [17]. In addition, the recombination probability for electron-hole pairs before reaching the avalanche region may be higher at liquid nitrogen temperature. Similar findings were observed when researching the quantum efficiency of silicon thin film solar cells [18]. It is not expected that thermal expansion significantly impacts the effective fill factor because the thermal expansion coefficient for silicon at liquid nitrogen temperatures amounts to only a 0.02% change in area, which is too small to explain the observed decrease in PDE.
@@ 5 Conclusions
An experimental setup has been built at the University of Washington to characterize SiPMs at liquid nitrogen temperature to assess the efficiency of the liquid argon veto readout system to be deployed in LEGEND-1000. The SiPM test stand consists of an integrating sphere suspended above a liquid nitrogen bath in a dewar housed inside a dark box. The SiPM is in thermal contact with a copper cold finger that is partially submerged in liquid nitrogen and the integrating sphere is in a room-temperature thermal bath. Measurement of both a KETEK SiPM and Hamamatsu SiPM PDE at liquid nitrogen temperature suggests the liquid argon veto readout system has a lower efficiency compared to the expected PDE for room temperature operation. We expect the same behavior to persist at liquid argon temperatures because they are only $10$ K away from the measured liquid nitrogen temperatures: past studies have shown that the PDE at $87$ K for similar wavelengths is only $1.1$ times larger than the PDE at $77$ K [15]. The reduction in PDE seems to be caused by temperature dependencies to the intrinsic quantum efficiency and avalanche triggering probability of the SiPM. Changes to the effective fill factor may be present but are most likely negligible in comparison to the changes in quantum efficiency and avalanche triggering probability. The measurement of a decrease in PDE greater than 20% for all overvoltages at cryogenic temperatures for a SiPM sensitive to the visible spectrum is significant information for assessing and improving the background suppression capability of LEGEND-1000.
@@ Acknowledgments
The authors appreciate the technical assistance and advice of J.F. Amsbaugh, T.H. Burritt, G. Giovanetti, G. Holman, A. Hostiuc, C.J. Nave, D. Peterson, G. Song, T.D. Van Wechel, I. Wang, and C. Wiseman.
@@ Appendix A Dark Characterization
Figure 7: Dark characterization data for the KETEK and Hamamatsu SiPMs at room and liquid nitrogen temperatures as a function of overvoltage. Room temperature data for the Ketek were taken with a $0.1$ V offset from the breakdown voltage, so liquid nitrogen data were acquired with the same offset. In addition, the number of events for the Ketek SiPM at $0.4$ V overvoltage at liquid nitrogen temperature was too low to analyze due to the extremely low dark count rate, so the data point was omitted.
The direct cross-talk was also measured at a variety of overvoltages for both devices. The cross-talk probability is larger at cryogenic temperatures for both devices, but there is a cross-over point where it becomes equal to the room temperature cross-talk probability, as seen in Fig. 7(c) and Fig. 7(d). This is somewhat surprising as the cross-talk probability depends on the gain (via the number of secondary photons produced) and the avalanche triggering probability. The gain does not change significantly at cryogenic temperatures, but the avalanche triggering probability might, as was discussed in Section 4 [15]. The difference between room and liquid nitrogen temperature cross-talk probabilities is also not as substantial as it is for the afterpulsing.
The Hamamatsu and KETEK afterpulsing rates are significantly higher at liquid nitrogen temperatures, as shown in Fig. 7(e,f). A higher afterpulsing rate is potentially explained by the increased charge trapping lifetime at cryogenic temperatures: trapped carriers are more likely to be re-released while the cell is recovering and are thus more likely to trigger an afterpulse [19]. In addition to charge trapping re-releases, afterpulsing can also occur due to secondary carriers drifting from the substrate into the recharging SPAD [20]. This type of afterpulsing could also increase due to increased mobility of charge carriers at cryogenic temperatures and longer SPAD recovery times: carriers in the bulk silicon are more easily drifted into the active region and trigger an afterpulse [21]. Fig. 2d shows that the higher afterpulsing rate degrades of the Hamamatsu charge spectrum at liquid nitrogen temperatures. A similarly high cryogenic afterpulsing rate for a Hamamatsu device was recently reported in [22].

### v2

@@ 4 Results
The relative PDE was measured at room temperature and at liquid nitrogen temperature for the KETEK SiPM and Hamamatsu SiPM, see Fig. 3. Errors on all three quantities in Eq. 2.2 were propagated during error calculations. In order to account for a ±$\pm 0.1$ V uncertainty on the breakdown voltage of the reference SiPM, a systematic uncertainty on $\mathrm{PDE}^{\mathrm{ref.}}$ of $1\%$ was found from the manufacturer’s datasheet and was used in the analysis. The room temperature was 72 F (295 K) held by an air conditioning unit and the liquid nitrogen temperature was measured to be 77 K with a calibrated PT1000 sensor placed in between the SiPM board and copper cold finger. These temperatures were stable within 1 K during data taking. The breakdown voltage of the KETEK SiPM was measured at room temperature to be 24.1 V; the breakdown voltage at 77 K was 20.6 V. The breakdown voltages were obtained by extrapolating the SiPM gain, $G$, as a function of reverse bias to $G=1$ and were confirmed with a reverse IV curve from a Keithley 2450 Sourcemeter®. Likewise, the breakdown voltage of the Hamamatsu SiPM was measured at room temperature to be 51.5 V; the breakdown voltage at 77 K was 42.2 V. Because the breakdown voltage depends on the temperature, the breakdown voltage was also used as a cross-check on the temperature. The breakdown voltages at liquid nitrogen temperature were determined by submerging the devices completely in liquid nitrogen in a separate experiment prior to the PDE measurement. While the reference SiPM was mounted on the lid, the breakdown voltage of the reference SiPM was found to not change after an LN fill, confirming that it was kept at room temperature.
Figure 3: Measurement of the relative PDE for a KETEK PM3325-WB-D0 SiPM and a Hamamatsu S13360-3050CS SiPM for a range of overvoltages. Measurements are done at room temperature and liquid nitrogen temperature using $562$ nm light. The datasheet values for the PDEs at the peak LED wavelength used are provided for reference. The error bands correspond to combined statistical and systematic (2% from the flux ratio $f$ and 1% from $\mathrm{PDE}^{\mathrm{ref.}}$) uncertainties.
The PDE was measured for a range of overvoltages and results show a significant decrease in PDE at liquid nitrogen temperature for all overvoltages. This decrease in PDE is present even when accounting for possible changes in the optical properties of the setup at different temperatures or between configurations. Any systematic impact of the light uniformity in the integrating sphere or SiPM alignment would be seen during room temperature PDE measurements. Our room temperature PDE measurements are in good agreement with the manufacturer’s datasheets, as seen in Fig. 3. Moreover, the agreement between manufacturer and measured room temperature PDE values indicates that all three SiPMs are consistent with their manufacturer reported values, within our experimental uncertainties; although, a common systematic offset between all three cannot be excluded. Any differences in SiPM alignment when the reference diode is moved to the lid for lid configuration PDE measurements are accounted for in measurements of the flux ratio, Eq. 2.5. Systematic studies were performed to determine if there was any temperature dependence in the transmission coefficient of the optical fibers connecting to the reference SiPM, the emission spectrum of the LED, or the reflectivity of the integrating sphere.
@@
Figure 4: The fluctuation in the effective PDE, $\frac{N_{\mathrm{SiPM}}^{\mathrm{lid}}}{N_{\mathrm{ref.}}^{\mathrm{lid}}}$, when only the optical fibers were cooled down sets a limit on the systematic error at $2\%$.
We also investigated the systematics arising from the finite width of the LED’s emission spectrum. Using Eq. 2.3 to calculate the reference SiPM PDE instead of using the reference SiPM’s PDE at the peak emission wavelength of the LED introduces only an average $0.02\%$ difference in our final cryogenic PDE results, which is below our statistical sensitivity. We therefore compute the cryogenic PDE using only the peak emission wavelength of the LED.
We checked that $\rho(\lambda)$ (the LED’s emission spectrum) and the integrating sphere’s optical properties were not functions of temperature. While the LED emission spectrum should not change because the LED is placed outside of the dewar, the optical fiber that connects the LED to the integrating sphere could have a changing transmission coefficient as a function of temperature. To constrain this possibility, we compared the number of photons detected at the reference SiPM at 77 K and 295 K while the test stand was in the lid PDE measurement configuration. The reference SiPM has a comparable spectral response to the SiPMs under test. The results, shown in Fig. 5, indicate that the reference SiPM detects roughly the same number of photons when the test stand is at both room and cryogenic temperatures. Thus there is neither a shift in the emission spectrum of the LED nor a shift in the reflectivity of the integrating sphere. This measurement also serves as an in situ cross-check that the optical fibers’ transmission probability is not temperature dependent.
We ensured that condensation was not forming inside the integrating sphere or on the surfaces of the SiPMs. First, prior to data taking, a liquid nitrogen fill was performed and the dewar insert was lifted to visually check that no condensation was present. Second, the number of photons detected by the reference SiPM at the lid does not change significantly between room temperature and cryogenic temperature measurements in Fig. 5: any change in reflectivity engendered by condensation inside the integrating sphere would cause a systematic change in the number of photons detected by the reference SiPM. The nitrogen gas boil-off from the liquid nitrogen ensures that condensation cannot form. Therefore, any changes in the optical properties of the test stand between the two temperatures do not account for the observed decrease in PDE.
@@
The decrease in PDE at 77 K indicates that some combination of the quantum efficiency, avalanche triggering probability, and effective fill factor may be changing with temperature. Similar reductions of the PDE at cryogenic temperatures have been found in other SiPMs at 10 K using a calibrated system [38], and at 77 K using uncalibrated systems [32, 33, 30, 39] and calibrated systems [29, 19]. These systems show similar decreases in PDE of roughly 30% at liquid nitrogen temperatures for wavelengths close to 560 nm. An FBK NUV-HD-cryo SiPM operated at 3 V overvoltage was measured to have a PDE drop from $\sim$24at room temperature to $\sim$18at 75 K for 525 nm light; however, the decrease in PDE at 5 V overvoltage is much smaller, dipping from $\sim$34down to only $\sim$29[32]. A VUV-sensitive Hamamtsu SiPM at 4 V overvoltage displayed a similar decrease in PDE from 28.4% at room temperature to 20.1% at 77 K for 570 nm light as reported in [29]. Because silicon is an indirect band-gap semiconductor and fewer phonons are present in the crystal lattice at liquid nitrogen temperature, the probability of charge carriers moving to the conduction band from the absorption of a photon will be lower. This is reflected in the increase of the absorption depth of photons at a given energy at liquid nitrogen temperatures; this effect could reduce the quantum efficiency [32, 33]. The band-gap of silicon also increases with decreasing temperature, but this may produce negligible effects on the quantum efficiency [40]. In addition, the recombination probability for electron-hole pairs before reaching the avalanche region may be higher at liquid nitrogen temperature. Similar findings were observed when researching the quantum efficiency of silicon thin film solar cells [41]. The avalanche triggering probability could also decrease due to carrier loss from carrier freeze-out [33]; however, the triggering probability may also increase slightly due to larger carrier mobility at liquid nitrogen temperatures. More work needs to be done to disentangle the many effects that might be causing the decrease in the PDE. It is not expected that thermal expansion significantly impacts the effective fill factor because the thermal expansion coefficient for silicon at liquid nitrogen temperatures amounts to only a 0.02% change in area, which is too small to explain the observed decrease in PDE.
@@ 5 Conclusions
An experimental setup has been built at the University of Washington to characterize SiPMs at liquid nitrogen temperature to assess the efficiency of the liquid argon veto readout system to be deployed in LEGEND-1000. The SiPM test stand consists of an integrating sphere suspended above a liquid nitrogen bath in a dewar housed inside a dark box. The SiPM is in thermal contact with a copper cold finger that is partially submerged in liquid nitrogen and the integrating sphere is in a room-temperature thermal bath. Numerous systematics studies were performed to ensure that no other optical properties of the experimental apparatus changed at liquid nitrogen temperatures. Measurement of both a KETEK PM3325-WB-D0 SiPM and Hamamatsu S13360-3050CS SiPM PDE at liquid nitrogen temperature suggests the liquid argon veto readout system has a lower efficiency compared to the expected PDE for room temperature operation. We expect the same behavior to persist at liquid argon temperatures because they are only $10$ K away from the measured liquid nitrogen temperatures: past studies have shown that the PDE at $87$ K for similar wavelengths is only $1.1$ times larger than the PDE at $77$ K [33]. The reduction in PDE seems to be caused by temperature dependencies to the intrinsic quantum efficiency and avalanche triggering probability of the SiPM. Changes to the effective fill factor may be present but are most likely negligible in comparison to the changes in quantum efficiency and avalanche triggering probability. Additional measurements of the PDE at different wavelengths at cryogenic temperatures should be pursued. The measurement of a decrease in PDE greater than 20% for all overvoltages at cryogenic temperatures for a SiPM sensitive to the visible spectrum is significant information for assessing and improving the background suppression capability of LEGEND-1000.
@@ Acknowledgments
This work was funded by the Department of Energy Office of Nuclear Science. The authors appreciate the technical assistance and advice of J.F. Amsbaugh, T.H. Burritt, G. Giovanetti, G. Holman, A. Hostiuc, C.J. Nave, D. Peterson, G. Song, T.D. Van Wechel, L. Varriano, I. Wang, and C. Wiseman. The authors would also like to thank A. de St. Croix for helpful discussion.
@@ Appendix A Dark Characterization
Figure 7: Dark characterization data for the KETEK PM3325-WB-D0 and Hamamatsu S13360-3050CS SiPMs at room and liquid nitrogen temperatures as a function of overvoltage. Room temperature data for the Ketek were taken with a $0.1$ V offset from the breakdown voltage, so liquid nitrogen data were acquired with the same offset. In addition, the number of events for the Ketek SiPM at $0.4$ V overvoltage at liquid nitrogen temperature was too low to analyze due to the extremely low dark count rate, so the data point was omitted.
The direct cross-talk was also measured at a variety of overvoltages for both devices. The cross-talk probability is larger at cryogenic temperatures for both devices, but there is a cross-over point where it becomes equal to the room temperature cross-talk probability, as seen in Fig. 7(c) and Fig. 7(d). This is somewhat surprising as the cross-talk probability depends on the gain (via the number of secondary photons produced) and the avalanche triggering probability. The gain does not change significantly at cryogenic temperatures, but the avalanche triggering probability might, as was discussed in Section 4 [33]. This could be evidence for an increased avalanche triggering probability arising from the larger mobility of charge carriers at liquid nitrogen temperatures [32]. The cross-over could indicate that the avalanche triggering probability saturates faster as a function of overvoltage in the liquid nitrogen case. The slight temperature dependence of the cross-talk probability agrees with the elevated cross-talk probability at liquid nitrogen temperatures reported in [42]. The difference between room and liquid nitrogen temperature cross-talk probabilities is also not as substantial as it is for the afterpulsing.
The Hamamatsu and KETEK afterpulsing rates are significantly higher at liquid nitrogen temperatures, as shown in Fig. 7(e,f). A higher afterpulsing rate is potentially explained by new charge traps or by the increased charge trapping lifetime at cryogenic temperatures: trapped carriers are more likely to be re-released while the cell is recovering and are thus more likely to trigger an afterpulse [43]. Carrier freeze-out could lead to the activation of new shallow charge traps that could explain both the high afterpulsing rate as well as the decrease in PDE via a decrease in the avalanche triggering probability due to carrier loss [33]. In addition to charge trapping re-releases, afterpulsing can also occur due to secondary carriers drifting from the substrate into the recharging SPAD [44]. This type of afterpulsing could also increase due to increased mobility of charge carriers at cryogenic temperatures and longer SPAD recovery times: carriers in the bulk silicon are more easily drifted into the active region and trigger an afterpulse [45]. Fig. 2d shows that the higher afterpulsing rate degrades the Hamamatsu charge spectrum at liquid nitrogen temperatures. A similarly high cryogenic afterpulsing rate for a Hamamatsu device was recently reported in [46].